Suppose real GDP is growing at 4 percent, the money supply is growing at 11 percent, the velocity of money is constant, and the real interest rate is 6 percent.

a. What is the current inflation rate aand nominal inerest rate?
b. If the money supply growth rate inceases to 15 percent, how will your answer in part (a) change?

c. If you were an investor, how would the change in the money supply growth affect your real p rae?trofitability, assuming that you now receive the new nominal inerest?
d. Based on your previous answers, would you prefer a fixed or a floating interest on your investment? Which you would prefer if you thought the money supply growth was going to be reduced?
1
Expert's answer
2012-09-06T11:56:19-0400
(a) M/P=Y/V
Y - growing by 4% so coefficient = 1.04
M - growing by 11% so coefficient = 1.11
V - constant =1
1.11/P=1.04/1
P=1.11/1.04≈1.067≈+6.7%

r-real interest rate
i-nominal interest rate
π-inflation
(1+r)=(1+i)/(1+π)
1.06=(1+i)/(1+0.067)
i=(1.06*1.067)-1
i≈0.131≈13.1%
______ ______ ______ ______

(b)
M/P=Y/V
Y - growing by 4% so coefficient = 1.04
M - growing by 15% so coefficient = 1.15
V - constant =1
1.15/P=1.04/1
P=1.15/1.04≈1.1058≈+10.58%

r-real interest rate
i-nominal interest rate
π-inflation
(1+r)=(1+i)/(1+π)
1.05=(1+i)/(1+0.1058)
i=(1.05*1.1058)-1
i≈0.1611≈16.11%
______ ______ ______ ______

i¹≈13.1%
i²≈16.11%
Δi=(16.11/13.1)-1≈
≈+22.98%

Nominal interest rate changed for ≈+22.98% or ≈+3.01 points.

(c)
My profitability will not change because the real interest rate didn't
change.

(d)
I prefer floating interest. If M is reduced, than I prefer fixed interest.